Find the equations of the tangent and normal to each of the following conics, the lengths of the subtangent and subnormal, then trace the curve showing these lines.


y = x2 – 6x + 4 at (4, -4).


1
Expert's answer
2021-11-30T13:07:14-0500
"y=x^2-6x+4"

"y'=2x-6"


Point "(4,-4)"

"slope=m_1=y'(4)=2(4)-6=2"


Tangent line is


"y-(-4)=2(x-4)"

The equation of the tangent line in slope-intercept form


"y=2x-12"



Find the equation of the normal at the point "(4, -4)"


"slope=m_2=-\dfrac{1}{m_1}=-\dfrac{1}{2}"

"y-(-4)=-\dfrac{1}{2}(x-4)"

The equation of the normal line in slope-intercept form


"y=-\dfrac{1}{2}x-2"

"TA"   is defined to be the subtangent at "P". "AN"  is called the subnormal.

The lengths "PT"  and "PN"  are called the tangent and normal.


"A(4, 0), T(6, 0), N(-4, 0)"


"PT=\sqrt{(6-4)^2+(0-(-4))^2}=2\sqrt{5}"

"PN=\sqrt{(-4-4)^2+(0-(-4))^2}=4\sqrt{5}"

"TA=2"

"AN=8"